if logx a, logx b, logx c, are in AP, prove it a, b, c are in GP
 

Grade:12th pass

2 Answers

Sudipan Datta
6 years ago
If log a,log b and log c are in A.P then,
Using property of AP,
2log b=log a+log c
or, log b^2 = log (ac)
or, b^2 = ac
The resultant expression is the general expression of G.P.
Therefore a,b,c are in G.P
Shailendra Kumar Sharma
6 years ago
Loga logb logc are in AP i.e. 2logb=loga + logcNow that gives logb^2=log(a*c)Thus takin both sides on x power gives b^2 =acThat proves a,b,c are in GP